# On weak solutions in Einstein theory and beyond

### v2

@@ I Introduction
It is widely believed that quantum gravity can resolve the singularities that are endemic to general relativity. However, what should replace the singularity in a gravitational collapse scenario is a matter of open debate Fazzini:2023ova ; Fazzini:2023scu . One interesting possibility is that the singularity is replaced by a bounce: matter falls inward, crosses an apparent horizon, thus forming a black hole, and then reaches a minimal radius, at which point it starts expanding and eventually exits the apparent horizon. The idea is inspired by loop quantum gravity (LQG), which predicts the geometry of spacetime to be discrete at a microscopic level and hence posits the existence of a minimal length. It has been concretely realized in phenomenological or effective models of gravitational collapse Husain:2022gwp ; Husain:2021ojz ; Cipriani:2024nhx ; Giesel:2023hys which incorporate LQG-inspired quantum corrections to classical models of inhomogeneous dust collapse in the framework of LTB spacetimes.
@@ II Weak solutions in dust collapse
The name stems from the fact that in any compact region of space $[r_{1},r_{2}]$,
[…]
Hyperbolic conservation laws have the property that solutions to smooth initial data become discontinuous in a finite time; in dust collapse, such discontinuities correspond to points where shells of dust have crossed each other Husain:2022gwp ; Fazzini:2023ova , the so-called shell-crossing singularities (SCS). Evolution of the system beyond such points can be studied using the integrated equation above. More precisely, suppose that there is a shell crossing singularity at $r=R(t)$, and suppose that $\lim_{r\to R^{\pm}}u=u^{\pm}$; then it can be shown leveque-92 using (9) that the discontinuity at $r=R(t)$ propagates according to the equation
@@ III Signature change and metric continuity in PG coordinates
It is immediate to notice that the induced metric as seen from both the interior and exterior changes signature during the evolution of the shell, and since ${\rm d}s^{2}$ gives the proper time interval of the shell (when $\theta$ and $\phi$ are fixed), this leads to a proper time interval that becomes null when $R=R_{S}/4$, and changes sign for $R<R_{S}/4$. As pointed out in Mars:1993mj ; Senovilla:2018hrw hypersurfaces that change signatures can arise in solutions of Einstein theory, but whether they provide nonphysical consequences or not has to be checked case by case. In our case, this implies that when the shell is at $R>R_{S}/4$ ($R<R_{S}/4$), then only a time-like (space-like) particle can co-move together with the shell surface, while when $R=R_{S}/4$ only a null particle can. Even if this kind of spacetime can be mathematically defined, it seems unrealistic from a physical perspective, since the shock is generated by a thin shell of dust which should move in a time-like fashion.
Since the only requirement to find the shock velocity (15) is continuity of the induced metric in PG coordinates — a plausible assumption at first sight—the question is how the dynamics can be nonphysical. The reason is subtle and concerns the behavior of the PG time coordinate across the shock. The discussion so far is based on the assumption that a single PG chart can be used both for the spacetime interior to the thin shell (which trivially gives the Minkowski metric), and for the exterior (that gives the Schwarzschild metric in PG coordinates). However, the time flowing in these two spacetimes, say $t_{+}$ for the exterior and $t_{-}$ for the interior, is extremely different; this is evident in the fact that a Schwarzschild metric in PG coordinates causes gravitational time dilation, whereas a Minkowski metric does not. When we require equality between (13) and (14), we require not only continuity of the induced metric on the shell, but also that ${\rm d}t_{-}^{2}={\rm d}t_{+}^{2}$, which means that the flow of PG time is the same as measured from both sides of the shell. This is a stronger assumption than mere continuity, since it does not account for a possible coordinate discontinuity of the PG time at the shock, and forces the signature of the shock induced metric to change during the evolution.
@@ IV PG time and Israel junction conditions
In light of the foregoing remarks, the only way to ensure a signature-preserving dynamics and metric continuity at the same time seems to be to take the interior and exterior PG times $t_{-}$ and $t_{+}$ to be different. A natural avenue to explore this point is to study the dynamics of thin shell using Israel junction conditions Poisson:2009pwt ; Israel:1966rt , and compare the outcome with the previous result. In this formalism, one starts by introducing separate coordinates on either side of the shell. Imposing metric continuity (the first junction condition) allow the Einstein equations to hold in a distributional sense, which give rise to jump relationships between the extrinsic curvature of the shell and its stress-energy tensor. Together, these conditions then serve to fix the dynamics of the shell completely. If this calculation is done in PG coordinates, assuming the coordinates $\{t_{\pm},r_{\pm},\theta,\phi\}$ on the two sides of the shell, one can see quite explicitly that (i) $t_{+}$ is not the same as $t_{-}$ and (ii) the shell always remains timelike. Since the latter fact gives us good reason to believe that the dynamics is physical, we are apt to take the equality of $t_{+}$ and $t_{-}$, which has to be assumed to get a weak solution with continuous induced metric, with a grain of salt. Thus, it will be instructive to run through the formalism of Israel junction conditions applied to a thin shell.
@@
Notice that by requiring ${\rm d}s^{2}|_{\theta,\phi=const.}=-{\rm d}\tau^{2}$ implies signature preservation of the induced metric at the shell. In other words, for metric continuity to be compatible with a signature-preserving metric, the PG times for the exterior and interior must be related to the shock velocities with respect to the exterior and interior by the above equations. However, this equation now does not fix the shock velocity. To fix the dynamics completely, one has to use the relation between the induced stress tensor and the extrinsic curvature of the shell to fix the dynamics. Since the expression for the extrinsic curvature of the shell is quite complicated in PG coordinates, we will instead write the interior and exterior metric in Minkowski and Schwarzschild coordinates respectively, and after having obtained an equation for the shell velocity, revert to PG coordinates to exhibit the relationship between the interior and exterior PG times. Thus, we write

### v3

@@ I Introduction
It is widely believed that quantum gravity can resolve the singularities that are endemic to general relativity. However, what should replace the singularity in a gravitational collapse scenario is a matter of open debate. One interesting possibility is that the singularity is replaced by a bounce: matter falls inward, crosses an apparent horizon, thus forming a black hole, and then reaches a minimal radius, at which point it starts expanding and eventually exits the apparent horizon. The idea is inspired by loop quantum gravity (LQG), which predicts the geometry of spacetime to be discrete at a microscopic level and hence posits the existence of a minimal length. It has been concretely realized in phenomenological or effective models of gravitational collapse Husain:2022gwp ; Bobula:2023kbo ; Lewandowski:2022zce ; Husain:2021ojz ; Cipriani:2024nhx ; Giesel:2023hys ; Alonso-Bardaji:2023qgu ; Han:2023wxg ; Cafaro:2024vrw which incorporate LQG-inspired quantum corrections to classical models of inhomogeneous dust collapse in the framework of LTB spacetimes.
@@ II Weak solutions in dust collapse
The name stems from the fact that in any compact region of space $r_{1},r_{2}]$,
$\frac{{\rm d}}{{\rm d}t}\int_{r_{1}}^{r_{2}}dr\,u(t,r)=f(r_{2},u(t,r_{2}))-f(r_{1},u(t,r_{1}))~,$ (9)
which means that the quantity $\int_{r_{1}}^{r_{2}}dr\,u(t,r)$ is conserved in the sense that it can change due only to the flux $f(r,u)$ through the boundaries $r_{1}$ and $r_{2}$.
Hyperbolic conservation laws have the property that solutions to smooth initial data become discontinuous in a finite time; in dust collapse, when such discontinuities arise in the matter region, they correspond to points where shells of dust have crossed each other Husain:2022gwp ; Fazzini:2023ova , the so-called (physical) shell-crossing singularities (SCS). Evolution of the system beyond such points can be studied using the integrated equation above. More precisely, suppose that there is a shell crossing singularity at $r=R(t)$, and suppose that $\lim_{r\to R^{\pm}}u=u^{\pm}$; then it can be shown leveque-92 using (9) that the discontinuity at $r=R(t)$ propagates according to the equation
@@ III Signature change and metric continuity in PG coordinates
It is immediate to notice that the induced metric as seen from both the interior and exterior changes signature during the evolution of the shell, and since ${\rm d}s^{2}$ gives the proper time interval of the shell (when $\theta$ and $\phi$ are fixed), this leads to a proper time interval that becomes null when $R=R_{S}/4$, and changes sign for $R<R_{S}/4$. As pointed out in Mars:1993mj ; Senovilla:2018hrw hypersurfaces that change signatures can arise in solutions of Einstein theory (or its distributional form), but whether they provide nonphysical consequences or not has to be checked case by case. In our case, this implies that when the shell is at $R>R_{S}/4$ ($R<R_{S}/4$), then only a time-like (space-like) particle can co-move together with the shell surface, while when $R=R_{S}/4$ only a null particle can. Even if this kind of spacetime can be mathematically defined, it seems unrealistic from a physical perspective, since the shock is generated by a thin shell of dust which should move in a time-like fashion.
Since the only requirement to find the shock velocity (15) is continuity of the induced metric in PG coordinates — a plausible assumption at first sight—the question is how the dynamics can be nonphysical. The reason is subtle and concerns the behavior of the PG time coordinate across the shock. The discussion so far is based on the assumption that a single PG chart can be used both for the spacetime interior to the thin shell (which trivially gives the Minkowski metric), and for the exterior (that gives the Schwarzschild metric in PG coordinates). However, the time flowing in these two spacetimes, say $t_{+}$ for the exterior and $t_{-}$ for the interior, is extremely different; this is evident in the fact that a Schwarzschild metric in PG coordinates causes gravitational time dilation, whereas a Minkowski metric does not. When we require equality between (13) and (14), we require not only continuity of the induced metric on the shell, but also that ${\rm d}t_{-}^{2}={\rm d}t_{+}^{2}$, which means that the flow of PG time is the same as measured from both sides of the shell. This is a stronger assumption than mere metric continuity, since it does not account for a possible coordinate discontinuity of the PG time at the shock, and forces the signature of the shock induced metric to change during the evolution.
@@ IV PG time and Israel junction conditions
In light of the foregoing remarks, the only possibility to ensure a signature-preserving dynamics and metric continuity at the same time seems to allow the interior and exterior PG times $t_{-}$ and $t_{+}$ to be different. A natural avenue to explore this point is to study the dynamics of thin shell using Israel junction conditions Poisson:2009pwt ; Israel:1966rt , and compare the outcome with the previous result. In this formalism, one starts by introducing separate coordinates on either side of the shell. Imposing metric continuity (the first junction condition) allows the Einstein equations to hold in a distributional sense, which gives rise to jump relationships between the extrinsic curvature of the shell and its stress-energy tensor. Together, these conditions then serve to fix the dynamics of the shell completely. If this calculation is done in PG coordinates, assuming the coordinates $\{t_{\pm},r_{\pm},\theta,\phi\}$ on the two sides of the shell, one can see quite explicitly that (i) $t_{+}$ is not the same as $t_{-}$ and (ii) the shell always remains timelike. Since the latter fact gives us good reason to believe that the dynamics is physical, we are apt to take the equality of $t_{+}$ and $t_{-}$, which has to be assumed to get a weak solution with continuous induced metric, with a grain of salt. Thus, it will be instructive to run through the formalism of Israel junction conditions applied to a thin shell.
@@
Notice that by requiring ${\rm d}s^{2}|_{\theta,\phi=const.}=-{\rm d}\tau^{2}$ implies signature preservation of the induced metric at the shell. In other words, for metric continuity to be compatible with a signature-preserving metric, the PG times for the exterior and interior must be related to the shock velocities with respect to the exterior and interior by the above equations. However, this equation now does not fix the shock velocity. To fix the dynamics completely, one has to use the relation between the induced stress tensor and the extrinsic curvature of the shell to fix the dynamics. Since the expression for the extrinsic curvature of the shell is quite complicated in PG coordinates, we will instead write the interior and exterior metric in Minkowski and Schwarzschild coordinates respectively, and after having obtained an equation for the shell velocity, revert to PG coordinates to exhibit the relation between the interior and exterior PG times. Thus, we write
@@ VI A different post-bounce dynamics for the shock wave model
Another approach to this problem, developed in the literature, deserves attention. It has been shown Friedman:1997gr ; Fiamberti:2007 that one can recover the classical Israel dynamics for the thin shell starting from a variational principle, using an action that includes boundary terms properly accounting for the thin-shell matter content. This approach has recently been employed to derive an effective thin-shell dynamics Sahlmann:2025fde . While the resulting dynamics reproduces the correct classical limit, it cannot be considered fully reliable, as it suffers from issues similar to those of the weak approach. In particular, it does not allow the imposition of the first Israel junction conditions (22), (23) in the deep quantum region, which in turn leads to a complex-valued lapse function. Even when a relaxed version of these conditions is imposed, the time-like behavior of the thin shell—which is guaranteed automatically in the Israel framework—is recovered only for small black hole masses. Future studies will aim to clarify why this approach works at the classical level but fails at the effective level.